The length of a rectangle road sign is 2 feet less than three times its width. Find the dimensions if the perimeter is 28 feet
Answer by satyareddy22(84)

Let length be l,breadth be b
perimeter = 28 feet
2(l+b)=28feet
=>l+b = 14----------(1)
3b-l = 2(given)-----(2)
solving (1) & (2)
l=10,b=4